OpenStudy (anonymous):

Choose the correct slope of the line that passes through the points (-4, 8) and (-3, -6).

OpenStudy (anonymous):

Find the gradient of the slope y2-y1/x2-x1 Then use the equation y-y1=m(x-x1)

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &({\color{red}{ -4}}\quad ,&{\color{blue}{ 8}})\quad &({\color{red}{ -3}}\quad ,&{\color{blue}{ -6}}) \end{array} \\\quad \\ slope = {\color{green}{ m}}= \cfrac{rise}{run} \implies \cfrac{{\color{blue}{ y_2}}-{\color{blue}{ y_1}}}{{\color{red}{ x_2}}-{\color{red}{ x_1}}}\)